### v1

# EEG Emotion Copilot: Pruning LLMs for Emotional EEG Interpretation with Assisted Medical Record Generation

Abstract
In the fields of affective computing (AC) and brain-machine interface (BMI), the analysis of physiological and behavioral signals to discern individual emotional states has emerged as a critical research frontier. While deep learning-based approaches have made notable strides in EEG emotion recognition, particularly in feature extraction and pattern recognition, significant challenges persist in achieving end-to-end emotion computation, including real-time processing, individual adaptation, and seamless user interaction. This paper presents the EEG Emotion Copilot, a system leveraging a lightweight large language model (LLM) operating in a local setting. The system is designed to first recognize emotional states directly from EEG signals, subsequently generate personalized diagnostic and treatment suggestions, and finally support the automation of electronic medical records. The proposed solution emphasizes both the accuracy of emotion recognition and an enhanced user experience, facilitated by an intuitive interface for participant interaction. We further discuss the construction of the data framework, model pruning, training, and deployment strategies aimed at improving real-time performance and computational efficiency. Privacy concerns are also addressed, with a focus on ethical data collection, processing, and the protection of users’ personal information. Through these efforts, we aim to advance the application of AC in the medical domain, offering innovative approaches to mental health diagnostics and treatment.
Index Terms:
Lightweight LLM, Model pruning, Emotion recognition, Assisted electronic medical record
I Introduction
The application of affective computing (AC) in brain-machine interfaces (BMI) is emerging as a key research frontier. Affective computing [1] seeks to identify emotional states by analyzing physiological and behavioral signals, such as electroencephalograms (EEG) [2] [3], heart rate [4], facial expressions [5], and voice [6]. Advances in this field have expanded the possibilities for BMI technology, particularly in enhancing human-computer interaction [7], and increasing its potential for practical applications in medical rehabilitation [8] and other areas [9].
The powerful learning capabilities of deep neural networks (DNNs) have driven rapid advancements in AC for emotion recognition via EEG signals [10], with numerous deep learning approaches proposed [11], [12], [13]. However, achieving end-to-end EEG-based AC remains challenging. Large language models (LLMs) [14], [15] built on the Transformer architecture [16] have demonstrated strong contextual understanding [17]. Models such as EEG-GPT [18] and ChatGPT-BCI [19] focus primarily on brain state identification but have yet to effectively address diverse, human-centered tasks. The integration of LLMs with EEG signals to generate personalized diagnoses, treatment plans, and assist in electronic medical records remains an urgent and complex challenge with the following limitations.
Firstly, building a suitable human-based corpus from diverse databases for LLMs is a challenging task. While public datasets like Wikitext [20] and C4 [21], along with specialized datasets like SEED [22] and FACED [23], offer substantial prior knowledge, this information remains fragmented. The challenge lies in linking these datasets to create a unified data structure that facilitates language model development and enables accurate responses through appropriate prompts.
Secondly, EEG signals can be viewed as a form of long text data [24], yet the signal sequences from a single channel often exhibit significant redundancy for current [25], [26] and traditional machine learning approaches [27], [28]. LLMs require a considerable number of tokens [16] to process this redundancy, resulting in unnecessary computational overhead. Additionally, many AC tasks rely on data from 32, 64, or more channels, further diminishing computational efficiency and presenting challenges for real-time performance. Therefore, effective data compression is critical to optimizing lightweight LLMs for real-time applications.
Finally, lightweight LLMs are optimally run locally to safeguard participant privacy and data security. While tasks based on public LLMs, such as EEG-GPT [18], appear feasible, many industry models necessitate powerful machines. For specific local tasks, complex DNNs are often excessively redundant. Therefore, effective model pruning is essential to achieve a lightweight language model suitable for local execution.
Therefore, we propose EEG Emotion Copilot, an intelligent system based on a lightweight, locally-running language model that utilizes EEG signals to perform emotion status recognition and subsequently generate corresponding diagnosis and treatment plans. This aims to automatically create assisted electronic medical records, as illustrated in Figure 1. In the following sections, we will detail the construction of an effective data framework, the process of model pruning, model training, and final model deployment.
Figure 1: Flowchart of EEG Emotion Copilot. The original EEG signal is preprocessed and transformed via wavelet to shorten the signal length. The final prompt is constructed using the initialized prompt, demographic data, emotional label, and treatment as training data. Using the Qwen2-0.5B-pretrained model [29] as an example, we prune it by 50% and restore its performance with Wikitext [20], C4 [21], and other datasets to obtain the pruned Qwen2-0.5B model. During the Model Fine-tuning stage, a Warm-up phase gradually increases the learning rate, and Lora [30] is used for fine-tuning. Finally, the model is deployed, utilizing the RAG (Retriever-Reader-Generator) technique [31] to enhance retrieval performance, and a dialogue method in the graphical interface to improve interactivity.
II RELATED WORK
II-A Data Structure in Prompting LLMs
For specific tasks in LLM prompt engineering, tailored prompts are essential to obtain satisfactory answers [32]. Although datasets like Wikitext [20] and C4 [21] contain extensive knowledge, constructing specific datasets is necessary for optimal performance in particular tasks. For participation-centered tasks, incorporating more personal information can yield better answers. In much of the work on AC, researchers focus on distinguishing emotions using convolutional neural networks (CNNs) or recurrent neural networks (RNNs) [33], [34], with EEG signals primarily used to assess model quality [35], [36], [37], [38], [39]. The excessive length of input EEG data is often overlooked, leading to low memory and computational costs. However, in practical deployment scenarios, the EEG data length significantly impacts reasoning effectiveness [40], [41]. LLMs must split text sequences into tokens for processing, with computational and memory requirements increasing exponentially with input length. This poses challenges for LLMs deployment, especially on low-resource devices or in real-time scenarios. Although data chunking strategies [42] can split long sequences for training and deployment, our experiments show limited effectiveness in practice. Segmentation increases data volume and complexity, complicating category identification. Therefore, a reasonable data compression process is necessary to LLMs.
II-B Model Pruning in LLMs
Model pruning [43] has been widely applied in DNNs to reduce computational complexity and mitigate model hallucinations [44]. Frankle and Carbin’s lottery ticket hypothesis [45] posits that small subnetworks, when trained from scratch, can achieve comparable performance. Li et al. [46] and Han et al. [47] introduced filter and weight pruning techniques, while Michel et al. [48] explored attention head pruning in Transformer models. Zhuang et al. [49] employed gradient-based structured pruning, and Liu et al. [50] proposed MetaPrune, which combines pruning with neural architecture search (NAS). Guo et al. [51] focused on channel pruning for efficient inference. Comprehensive surveys by Blalock et al. [52] and He et al. [53] have highlighted the effectiveness of various pruning techniques and outlined future research directions. In the context of LLMs, methods such as SparseGPT [54] and LLM-Pruner [55], primarily based on the LLaMA model [56], have demonstrated promising results on public datasets and claim faster performance on local machines. However, the hardware used for testing, such as RTX 4090 GPUs, significantly outperforms standard PCs, leaving the inference cost still relatively high. Thus, for local computations with limited resources, there remains a need for more efficient and feasible pruning methods to further reduce computational costs.
II-C Assisted Electronic Medical Record Generation
Electronic medical records (EMRs) date back to the 1960s [57]. With advancements in computer science and artificial intelligence (AI), EMRs have evolved from merely storing and managing patient information to incorporating more complex functionalities. Today, AI models are increasingly integrated into diagnosing mental illnesses and detecting internal organ pathologies, enhancing the capabilities of EMRs. These models can be embedded within ”assisted electronic medical record” shown in Figure 2, where multiple copilots collaborate to process diagnostic data and produce accurate results. For simpler tasks, a single copilot may suffice. In response, we developed a new EEG Emotion Copilot that leverages a lightweight local LLM to perform emotion recognition based on EEG signals and provide corresponding diagnosis and treatment plans.
Figure 2: Schematic diagram of the assisted electronic medical record generated by intelligent copilots. It includes Basic information, Medical History, Physical Examination Records, Diagnosis Information, Treatment Plan, Laboratory Results, Follow-up Records, and Medical Expenses. Multiple copilots diagnose symptoms in corresponding fields and give comprehensive diagnosis results combined with laboratory results, and provide treatment plans.
III METHOD
III-A Data Structures of Prompt in EEG Emotion Copilot
Fine-tuned LLMs often achieve strong performance and generate end-to-end outputs, particularly for simple classification tasks. For example, distinguishing whether a subject’s emotional state is positive or negative based on EEG data is relatively straightforward. However, distinguishing more nuanced emotions, such as those in the nine-class FACED dataset [23], or accurately identifying emotions in real-world scenarios, remains challenging. To address this, we incorporate demographic features, including gender, age, and facial characteristics, into the data. General LLMs are typically trained on broad datasets like Wikitext [20] and Common Crawl [21] for general knowledge and logical reasoning. For domain-specific applications, they are trained on specialized datasets, such as BioBERT [58] for biomedicine, ClinicalBERT [59] for medical texts, and FinBERT [60] for financial texts. These datasets often consist of declarative sentences, but for tasks like logical reasoning, data are stored as question-answer pairs, as shown in Figure 3. For EEG-based emotion diagnosis and treatment planning, a question-answer system offers a more effective solution.
Figure 3: Comparison of three data structures: the general data structure (left), the question-answering system data structure (middle), and the data structure used in the proposed EEG Emotion Copilot (right).
In the data structure of the proposed EEG Emotion Copilot, we initialize the system with the prompt showed in the right part of Figure 2: ”You are an EEG emotion analyzer. I will input the patient’s personal information and EEG signals from specific electrode positions. Please infer the patient’s current emotional state and provide a detailed diagnosis along with personalized treatments based on your knowledge base.” Following initialization, we input the subject’s demographic data, including gender, age, and facial features, to assess their emotional state. Next, the acquired EEG signals are processed. As illustrated in Figure 1, preprocessing steps such as artifact removal and re-referencing are applied to the signals, followed by wavelet transformation for signal compression. These processed signals are then used as input. Finally, the structure is completed by adding emotion labels and corresponding treatment plans in the second and third sections, respectively.
III-B Model Pruning in EEG Emotion Copilot
Although some pruning schemes may achieve better results, we use torch pruning [61] in this paper for its flexibility and versatility. We set the pruning_ratio to 0.5, reducing the model’s parameters by half. To restore performance, we train and fine-tune the pruned model on a public dataset. We then pseudo-quantize the model weights by pseudo-quantizing [62] the top 1% of the most important weights, identified by their calculated “importance.” The remaining weights are left unchanged. For details, please see Figure 4.
Figure 4: Flowchart of model pruning in EEG Emotion Copilot: The upper left presents a rough framework diagram, while the right side illustrates torch pruning with a 0.5 pruning ratio. In this process, the embed_tokens in the embedding layer is reduced by half, and the input_feature and out_feature of Qwen2SdpaAttention and Qwen2MLP are halved compared to the original model. The final lm_head output remains unchanged, as it is directly tied to the vocabulary size. The lower left depicts a simplified torch pruning process.
Figure 5: Flowchart of electronic medical record generation implemented in EEG Emotion Copilot.
III-C Model Fine-tuning in EEG Emotion Copilot
Even when a pre-trained model is fine-tuned on a public dataset post-pruning, certain tasks may still require fine-tuning with a proprietary dataset. In this study, the Low-Rank Adaptation (LoRA) [30] has been employed, which introduces low-rank matrices in specific layers to adjust parameters, avoiding the need for complete retraining of all weights, thus enabling faster and more efficient adaptability. Additionally, a warm-up strategy is often applied at the onset of training to facilitate convergence and improve overall model performance.
III-D Model Deployment in EEG Emotion Copilot
Given the recent release of various model deployment tools, such as llama.cpp, which can quantify the model based on the specific scenario and convert it into gguf format, it can be directly utilized through ollama, llmstudio, or other tools. This process ultimately results in the model deployment part illustrated in Figure 1. Consequently, we will not delve into the details of model deployment and will proceed directly to the model application section.
III-E Assisted Electronic Medical Records
Different models need to be trained for various tasks. For emotion recognition, diagnosis, and treatment plan design, it is often sufficient to embed demographic data of different cases under various emotions along with the corresponding compressed EEG signals in the training data. In downstream applications, the collected participant data can be input into the fine-tuned model in a specific format to obtain the corresponding diagnosis results and treatment plans, as shown in Figure 5.
IV EXPERIMENTS
IV-A Implementation Details
Dataset: We utilize the FACED dataset [24] to construct our training dataset. This dataset includes recordings from 123 subjects, each with 32 EEG channels, as they watched 28 video clips, each lasting 30 seconds. These clips encompass nine emotions: Anger, Disgust, Fear, Sadness, Neutral, Amusement, Inspiration, Joy, and Tenderness. Each positive and negative emotion is represented by three clips, while the neutral emotion is depicted in four clips.
Training Details: In the model pruning phase, we set the pruning ratio to 0.5 and trained on the public dataset with a learning rate of 1e-5, incorporating L1 and L2 regularization terms. The training was conducted for 100 iterations. During the Lora fine-tuning phase, we maintained a learning rate of 1e-5, applied weight decay and gradient decay, and set LoRA_r to 8, LoRA_alpha to 32, and LoRA_dropout to 0.5. The fine-tuning process spanned 15 epochs, divided into 5 steps. Model pruning was executed on an A800 80G, while model training and fine-tuning were completed on a 4090 24G.
IV-B Performance Validation of Varied EEG Signal Length
Regarding the compressed signal length, we first constructed the signal in the FACED data and the corresponding demography according to the Data Structure in Figure 1. We implemented two types of compression methods and an original control group. The question-answering process and the corresponding results are shown in Figure 6. Although the Euclidean distance and Cosine similarity of Ans1 and Ans2 are similar to the Baseline for the same question, Ans1 is more consistent with the Baseline than Ans2, as Ans2 differs from the Baseline in terms of emotion recognition. Therefore, with limited resources, it is often more effective to compress the signal to a fixed length using wavelet transforms rather than a segmented approach.
IV-C Model Retraining Strategy Selection After Pruning
We explored two sequential retraining strategies to restore model performance after applying torch pruning, as illustrated in Figure 7. It is reasonable to first train the model on public datasets and then on a specific dataset, as public datasets are rich in categories and provide a comprehensive feature representation for the model’s learning. This approach allows the model to leverage the diversity of public datasets to learn more generalized features initially. Consequently, when the model encounters a specific dataset, it can quickly adjust and optimize its weights to meet the specific task requirements.
Alternatively, we first trained on a specific dataset and then fine-tuned on a public dataset. This method allows the model to initially focus on the features of a specific task and subsequently improve its generalization ability and robustness through exposure to a public dataset. However, to save computing resources, training solely on a specific dataset may suffice.
We conducted an experiment for validation, as shown in Figure 8, dividing the training process into five steps, each with three epochs. In the first four steps, we used 250Hz data from 55 subjects in the FACED dataset. We observed that the model’s answers in step 3 were more satisfactory than in step 4, despite minimal differences in training_loss, indicating that excessive training may be unnecessary or even detrimental for lightweight models. In step 5, we used 1000Hz data from 68 subjects in the FACED dataset, with the pre-trained from step 4. The model trained on a dataset with a similar structure but different information density provided richer answers than in step 4, confirming that the model had learned the data structure and corresponding answers. This validates our hypothesis that the pruned model can be effectively retrained on a specific dataset to provide specific answers.
IV-D Assisted Electronic Medical Record Generation with Our Copilot
Figure 9 illustrates an example of using our EEG Emotion Copilot to create an electronic medical record. Initially, the starting sentence of the prompt is provided, followed by the addition of demographic information of the subjects, including age, gender, and certain facial features. Finally, the preprocessed and compressed signal is incorporated to complete the prompt, which is then input into EEG Emotion Copilot. The EEG Emotion Copilot first analyzes the data to generate an answer, which is subsequently transformed into the record through format conversion.
Figure 6: Two demos of paired question and answer with signal lengths of 50 (model 1) and 500 (model 2) time points. The doctor first asks a question to the trained model 1 or model 2. For model 1, the wavelet transform compresses the original signal of each channel into 50 points as training data. For model 2, the wavelet transform initially compresses the original signal of each channel into 500 points, then divides the compressed channel signal into 10 segments. These ten sequential group data serve as training data for model 2. Ans1 and Ans2 are the answers from model 1 and model 2, respectively, while the baseline represents the real answer to the question.
Figure 7: Two types of retraining strategies. The first strategy involves initially training the model on a specific dataset tailored to the current task, followed by training on a public dataset. The second strategy begins with training the model on a public dataset, and subsequently fine-tuning it on a specific dataset.
Figure 8: Comparative experimental results of the model’s training process. The training is divided into five steps, each consisting of three epochs. The first four steps utilize 250Hz data from 55 subjects. Comparing the model’s answers in steps 3 and 4 reveals that the answer in step 3 is richer in content, with no significant difference in training loss between the two steps. This suggests that overtraining lightweight models may be redundant or even detrimental. In step 5, 1000Hz data from 68 subjects is used for training, employing the pre-trained model from step 4.
Figure 9: An example using our EEG Emotion Copilot to create the assisted electronic medical record. Initially, the prompt is provided as the starting sentence, followed by the addition of the subject’s demographic information, including gender, age, and certain facial features (female, 23 years old, negative). Finally, the preprocessed and compressed EEG signal is incorporated to complete the prompt, which is then input into EEG Emotion Copilot.
V Discussion and Conclusion
In this paper, we present EEG Emotion Copilot, a lightweight and locally-running LLM designed for emotional recognition using EEG signals and the generation of corresponding treatment plans for assisted electronic medical records. Our approach addresses several key challenges in the fields of AC and BMIs.
Firstly, we developed a comprehensive human-based dataset that integrates public knowledge datasets and specialized emotion datasets, enabling the construction of a robust data structure to drive the development of LLMs. By leveraging appropriate prompting strategies, we enhanced the model’s ability to provide contextually relevant responses.
Secondly, we addressed the issue of data redundancy inherent in EEG signal processing. Through efficient data compression techniques, we optimized the handling of long sequences of EEG data, significantly improving computational efficiency and facilitating real-time emotion computation.
Lastly, we emphasized the importance of patient privacy by ensuring that our model can run locally. We explored model pruning strategies to create a lightweight version of the language model, making it feasible to deploy in environments with limited computational resources while maintaining performance.
However, simply compressing signal data is computationally intensive in many cases, and in complex scenarios, more channel signals are needed. Therefore, if LLMs can generate dense channel signals using only a few channel data and then analyze the emotions of the subjects [63], [64], computational efficiency will be greatly improved.
Our findings indicate that EEG Emotion Copilot has the potential to revolutionize the recognition and treatment of emotions in clinical settings, ultimately enhancing human-computer interaction and improving patient outcomes. Future work will combine lightweight LLMs with multimodal data such as human facial expressions, speech, gestures to achieve greater accuracy in affective computing.