### v2

# EEG Emotion Copilot: Optimizing Lightweight LLMs for Emotional EEG Interpretation with Assisted Medical Record Generation

Abstract
In the fields of affective computing (AC) and brain-machine interface (BMI), the analysis of physiological and behavioral signals to discern individual emotional states has emerged as a critical research frontier. While deep learning-based approaches have made notable strides in EEG emotion recognition, particularly in feature extraction and pattern recognition, significant challenges persist in achieving end-to-end emotion computation, including real-time processing, individual adaptation, and seamless user interaction. This paper presents the EEG Emotion Copilot, a system optimizing a lightweight large language model (LLM) with 0.5B parameters operating in a local setting, which first recognizes emotional states directly from EEG signals, subsequently generates personalized diagnostic and treatment suggestions, and finally supports the automation of assisted electronic medical records. Specifically, we demonstrate the critical techniques in the novel data structure of prompt, model pruning and fine-tuning training, and deployment strategies aiming at improving real-time performance and computational efficiency. Extensive experiments show that our optimized lightweight LLM-based copilot achieves an enhanced intuitive interface for participant interaction, superior accuracy of emotion recognition and assisted electronic medical records generation, in comparison to such models with similar scale parameters or large-scale parameters such as 1.5B, 1.8B, 3B and 7B. In summary, through these efforts, the proposed copilot is expected to advance the application of AC in the medical domain, offering innovative solution to mental health monitoring. The codes will be released at https://github.com/NZWANG/EEG_Emotion_Copilot.
Index Terms:
Lightweight LLM, Model pruning, Model fine-tuning, Emotion recognition, Assisted electronic medical record
I Introduction
THE application of affective computing (AC) in brain-machine interfaces (BMI) is emerging as a key research frontier. Affective computing [1] seeks to identify emotional states by analyzing physiological and behavioral signals, such as electroencephalograms (EEG) [2] [3], heart rate [4], facial expressions [5], voice [6], etc. Advances in this field have expanded the possibilities for BMI technology, particularly in enhancing human-machine interaction [7], [8], and increasing its potential for practical applications in rehabilitation [9], [10] and other areas [11].
The powerful capabilities of deep neural networks (DNNs) have driven rapid advancements in AC for emotion recognition via EEG signals [12], [13], with numerous deep learning approaches proposed [14], [15]. Recently, large language models (LLMs) [16], [17] built on Transformer architecture [18] have demonstrated strong contextual understanding [19]. Models such as EEG-GPT [20] and ChatGPT-BCI [21] focus primarily on brain state identification but have yet to effectively address diverse, human-centered tasks. The integration of LLMs with EEG signals to generate personalized diagnoses, treatment plans, and assisted electronic medical records remains an urgent and complex challenge with the following limitations.
Figure 1: Flowchart of EEG Emotion Copilot. The original EEG signal is preprocessed and transformed via wavelet to shorten the signal length. The final prompt is constructed using the initialized prompt, demographic data, emotional label, and treatment as training data. Using the Qwen2-0.5B-pretrained model [22] as an example, we prune it with the radio of 50% and fine-tuning the pruned model directly with specific dataset. During the Model Fine-tuning stage, a warm-up phase gradually increases the learning rate, and LoRA [23] is used for fine-tuning. Finally, the model is deployed, utilizing the RAG (Retriever-Reader-Generator) technique [24] to enhance retrieval performance, and a dialogue deployment of the graphical interface to improve interactivity.
Firstly, building a suitable human-based multimodal corpus from diverse databases for LLMs is a challenging task. While public datasets like Wikitext [25] and Common Crawl [26], along with specialized datasets like SEED [27] and FACED [28], offer substantial prior knowledge, this information remains fragmented. The challenge lies in linking these datasets to create a unified data structure that facilitates language model development and enables accurate responses through appropriate prompts.
Secondly, EEG signals can be viewed as a form of long text data [29], yet the signal sequences from a single channel often exhibit significant redundancy for machine learning approaches [30], [31] [32], [33]. LLMs require a considerable number of tokens [18] to process this redundancy, resulting in unnecessary computational overhead. Additionally, many AC tasks rely on EEG data from 32, 64, or more channels, further diminishing computational efficiency and presenting challenges for real-time performance. Therefore, effective data compression is critical to optimizing lightweight LLMs for real-time applications.
Finally, lightweight LLMs are optimally run locally to safeguard participant privacy and data security. While tasks based on public LLMs, such as EEG-GPT [20], appear feasible, many industry models necessitate powerful machines. For specific local tasks, complex DNNs are often excessively redundant. Therefore, effective model pruning is essential to achieve a lightweight language model suitable for local execution.
Therefore, we propose EEG Emotion Copilot, an intelligent system based on a lightweight, locally-running language model that utilizes EEG signals to first perform emotion recognition, subsequently generate corresponding diagnosis and treatment plans, and finally create assisted electronic medical records, as illustrated in Fig. 1.
II RELATED WORK
II-A Lengthy EEG Signal Embedded in Prompt
For specific tasks in LLM prompt engineering, tailored prompts are essential to obtain reasonable answers [34]. Although datasets like Wikitext [25] and Common Crawl [26] contain extensive knowledge, constructing specific datasets is necessary for optimal performance in particular tasks. In most studies on AC, researchers focus on distinguishing emotions using convolutional neural networks (CNNs) or recurrent neural networks (RNNs) [35], with EEG signals primarily used to assess model quality [36], [37]. Recently, LLMs should split EEG sequences into tokens for processing, with computational and memory requirements increasing exponentially with input length [38]. This poses challenges for LLMs deployment, especially on low-resource devices or in real-time scenarios. Although data chunking strategies [39] can split long sequences for training and deployment, the EEG data length significantly impacts reasoning effectiveness [40], [41]. Therefore, a reasonable data compression method is necessary to LLMs.
II-B Model Pruning in Deep Learning
The Frankle and Carbin’s lottery ticket hypothesis [42] posits that small subnetworks, when trained from scratch, can achieve comparable performance. Based on it, model pruning [43] has been widely applied in DNNs to reduce computational complexity and mitigate model hallucinations [44]. For instance, Li et al. [45] and Han et al. [46] introduced filter and weight pruning techniques, while Michel et al. [47] explored attention head pruning in Transformer models. Zhuang et al. [48] employed gradient-based structured pruning, and Liu et al. [49] proposed MetaPrune, which combines pruning with neural architecture search (NAS). Guo et al. [50] focused on channel pruning for efficient inference. Comprehensive surveys by Blalock et al. [51] and He et al. [52] have highlighted the effectiveness of various pruning techniques and outlined future research directions. In the context of LLMs, methods such as SparseGPT [53] and LLM-Pruner [54], primarily based on the LLaMA model [55], have demonstrated promising results on public datasets and claim faster performance on local machines. However, the hardware used for inference, such as RTX 4090 GPUs, significantly outperforms standard PCs, leaving the inference cost still relatively high. Thus, for local computations with limited resources, there remains a need for more efficient and feasible pruning methods to further reduce computational costs.
II-C Assisted Electronic Medical Records
Electronic medical records (EMRs) date back to the 1960s [56]. With advancements in computer science and artificial intelligence (AI), EMRs have evolved from merely storing and managing patient information to incorporating more complex functionalities. Currently AI models are increasingly integrated into diagnosing mental illnesses and detecting internal organ pathologies, enhancing the capabilities of assisted EMRs. The sample of ”assisted electronic medical record” is shown in Fig. 2, where the intelligent copilot can process the multimodal diagnostic data and produce accurate results. In this study, we developed a new EEG Emotion Copilot that leverages a lightweight local LLM to perform emotion recognition based on EEG signals, and to further generate the corresponding assisted EMR.
Figure 2: Schematic diagram of the assisted electronic medical record generated by intelligent copilot. It can include Basic Information, Medical History, Physical Examination Records, Diagnosis Information, Treatment Plan, Laboratory Results, Follow-up Records, and Medical Expenses.
III METHOD
III-A Data Structures of Prompt in EEG Emotion Copilot
Fine-tuned LLMs often achieve strong performance to generate end-to-end outputs, particularly for simple classification tasks. For instance, distinguishing whether a subject’s emotional state is positive or negative based on EEG data is relatively straightforward. However, distinguishing more nuanced emotions, such as nine-class emotions in FACED dataset [28], or accurately identifying emotions in real-world scenarios, remains challenging.
For Wikitext or similar datasets, these datasets often consist of declarative sentences (left part of Fig. 3), but for tasks like logical reasoning, data are stored as question-answer pairs, as shown in the middle part of Fig. 3. Despite these advancements, current data structures encounter significant limitations when applied to EEG-based emotion diagnosis and treatment planning. These challenges arise particularly in integrating multimodal data and addressing task-specific requirements. To address these gaps, we propose a novel structure leveraging a question-answer system tailored to EEG-related tasks.
Figure 3: Comparison of three data structures as to prompt in LLMs: the general data structure (left), the question-answering system data structure (middle), and the data structure used in the proposed EEG Emotion Copilot (right).
In the data structure of the proposed EEG Emotion Copilot, we initialize the system with the prompt showed in the right part of Fig. 3: ”You are an EEG emotion analyzer. I will input the patient’s personal information and EEG signals from specific electrode positions. Please infer the patient’s current emotional state and provide a detailed diagnosis along with personalized treatments based on your knowledge base.” Following initialization, we input the subject’s demographic data, including gender, age, and facial features, to assess their emotional state. Next, the acquired EEG signals are processed. As illustrated in Fig. 1, preprocessing steps such as artifact removal and re-referencing are applied to the signals, followed by wavelet transformation for signal compression. These processed signals are then used as input. Finally, the structure is completed by adding emotion labels and corresponding treatment plans in the second and third sections, respectively.
III-B Model Pruning in EEG Emotion Copilot
In this study, for the flexibility and versatility, we applied the torch pruning [57] to perform the model pruning task with the pruning_ratio equal to 0.5, and the detailed process is showed in Fig. 4.
Firstly, based on the architecture of the Qwen2-0.5B model, the total number of layers $L$ is determined as follows:
$\mathcal{L}=Num(Qwen2\_layer)=Num(Embedding\_layer)$ (1)
$+Num(Qwen2Decoder\_layer)+Num(Lm\_head)=25.$
The input and output of each layer are represented as $f_{i}^{-}$ and $f_{j}^{+}$, respectively, where $0\leq i\leq 25,0\leq j\leq 25$ and $i,j\in\mathbb{N}^{+}$.
Next, we construct a layer dependency graph $\mathcal{F}$ according to [57]. Based on the symmetric mechanism of input $f_{i}^{-}$ and output $f_{j}^{+}$, the dependency graph $\mathcal{F}$ is defined as a $\mathcal{L}\times 2\mathcal{L}$ symmetric matrix, specifically as follows:
$\mathcal{F}\left(f_{i}^{-},f_{j}^{+}\right)$ $=\mathcal{F}\left(f_{j}^{+},f_{i}^{-}\right)$ (2)
$=$ $\begin{cases}1,&\text{if }f_{i}^{-}\leftrightarrow f_{j}^{+}\text{ or }\\ &\;\left(i=j\text{ and }\operatorname{sch}\left(f_{i}^{-}\right)=\operatorname{sch}\left(f_{j}^{+}\right)\right),\\ 0,&\text{otherwise}\end{cases}$
where $\text{sch}(p)$ operation is used to determine which parameters in the parameter group $p$ need to be retained. Formula 2 returns a boolean value, where $\text{sch}(f_{i}^{+})=\text{sch}(f_{j}^{-})$ indicates that the two parameter groups are subject to the same pruning scheme.
In addition, we use the $\mathcal{L}_{2}$ norm to calculate the importance of the parameter group corresponding to each layer [57]. Based on the Qwen2 architecture, we calculate the importance $\mathcal{I}$ of the Embedding layer, Self-Attention layer, MLP layer, and Linear layer in turn as follows:
$\small\mathcal{I}(\text{embed\_tokens})=\sum_{w\in\text{embed\_tokens}}\|w\|_{2}^{2},$ (3)
$\footnotesize\mathcal{I}(s_{i})=\sum_{w\in s_{i}}\|w\|_{2}^{2},\;S=\{s_{i}\}=\{q_{\text{proj}},k_{\text{proj}},v_{\text{proj}},o_{\text{proj}}\},$ (4)
$\footnotesize\mathcal{I}(m_{i})=\sum_{w\in m_{i}}\|w\|_{2}^{2},\;M=\{m_{i}\}=\{gate_{\text{proj}},up_{\text{proj}},down_{\text{proj}}\},$ (5)
$\small\mathcal{I}(\text{lm\_head})=\sum_{w\in\text{lm\_head}}\|w\|_{2}^{2}.$ (6)
Based on the above formula, we first group the parameters $\mathcal{P}$ to get set $\mathcal{P^{\prime}}$, as shown in Algorithm 1. Then, we proceed to prune the model, as shown in Algorithm 2.
Figure 4: Flowchart of model pruning in EEG Emotion Copilot: The upper left presents a rough framework diagram, while the right side illustrates torch pruning with a 0.5 pruning ratio. In this process, the input_feature and out_feature of Qwen2SdpaAttention and Qwen2MLP are halved compared to the original model. The final output of lm_head remains unchanged, as it is directly tied to the vocabulary size. The lower left depicts a simplified torch pruning process.
Algorithm 1 Parameter Grouping
1:Input: Dependency graph $\mathcal{F}$, parameter group set $\mathcal{P}$
2:Output: Group set $\mathcal{P}^{\prime}$
3:Initialize $\mathcal{P}^{\prime}\leftarrow\emptyset$
4:for all $i$in $\mathcal{L}$  do
5:     Create new group $p\leftarrow\{k\}$
6:     Expand group $p$:
7:     while there exists an unvisited node $j$ such that $\mathcal{F}$$(f_{i}^{-},f_{j}^{+})=1$ do
8:         $p\leftarrow p\cup\{j\}$
9:         Mark $j$ as visited
10:     end while
11:     $\mathcal{P}^{\prime}\leftarrow\mathcal{P}^{\prime}\cup p$
12:end for
13:return $\mathcal{P}^{\prime}$
III-C Model Fine-tuning in EEG Emotion Copilot
To restore the performance of the pruned model $\mathcal{M^{\prime}}$, we explored two fine-tuning strategies after applying torch-pruning [57], as illustrated in Fig. 5. Strategy 1 is only training the pruned model on the specific EEG dataset to focus the model on task-relevant features with Low-Rank Adaptation (LoRA) [23]. Strategy 2 firstly involves full fine-tuning the pruned model on public datasets Wikitext [25] and Common Crawl [26], to leverage their diverse categories to establish generalized features, and subsequently fine-tuning on the specific EEG dataset with LoRA. Moreover, a warm-up strategy is applied at the beginning of training to facilitate convergence and improve overall model performance.
Specifically, the model’s fine-tuning process on the specific EEG dataset is illustrated in Fig. 6, dividing the training process into five steps, each with three epochs. In steps 1–4, we used 250 Hz data from 55 subjects of FACED [28]. Interestingly, the model’s performance in step 3 surpassed that in step 4, despite minimal differences in training_loss, suggesting that excessive training may be unnecessary and even detrimental for lightweight models. In step 5, 1000 Hz data from 68 FACED subjects pre-trained on step 4 yielded richer responses, confirming the model’s capability to learn the data structure and generate task-specific answers effectively.
III-D Model Deployment in EEG Emotion Copilot
Given the various model deployment tools, such as llama.cpp, which can quantify the model based on the specific scenario and convert it into gguf format, it can be directly utilized through ollama, llmstudio, gradio or other tools. This process with gradio ultimately results in the model deployment part illustrated in Fig. 1.
Figure 5: Two fine-tuning strategies. Strategy 1 involves initially training the model on a specific EEG dataset tailored to the current task. Strategy 2 begins with training the model on public datasets, and subsequently fine-tuning it on a specific EEG dataset.
Figure 6: Model’s fine-tuning process on the specific EEG dataset. This fine-tuning process is divided into five steps, each consisting of three epochs. The first four steps utilize 250 Hz data from 55 subjects. Comparing the model’s answers in steps 3 and 4 reveals that the answer in step 3 is richer in content, with no significant difference in training loss between the two steps. This suggests that overtraining lightweight models may be redundant or even detrimental. In step 5, 1000 Hz data from 68 subjects is used for training, employing the pre-trained model from step 4.
Algorithm 2 Pruning Process
1:Input: Initial model $\mathcal{M}$, Parameter group set $\mathcal{P^{\prime}}$, importance function $\mathcal{I}$, contraction strength $\alpha$
2:Output: Pruned model $\mathcal{M}^{\prime}$
3:Calculate group importance:
4:for all $p^{\prime}\in\mathcal{P^{\prime}}$ do
5:     $\mathcal{I}(p^{\prime})\leftarrow\sum_{w\in\mathcal{P^{\prime}}}\|w\|_{2}^{2}$ (See Formula 3, 4, 5, 6)
6:end for
7:    $\mathcal{P}^{\prime\prime}\leftarrow\text{sort}(\mathcal{P^{\prime}},\text{by }\mathcal{I})$
8:Select groups for pruning based on importance:
9:for all $p^{\prime\prime}\in\mathcal{P}^{\prime\prime}$ do
10:     if $\mathcal{I}(p^{\prime\prime})$leq $\theta$ is satisfied then #$\theta$ is the threshold
11:         $\mathcal{P^{\prime\prime}}\leftarrow\mathcal{P^{\prime\prime}}\setminus p^{\prime\prime}$
12:     end if
13:end for
14:return $\mathcal{M}^{\prime}$
III-E Generation of Assisted Electronic Medical Records
Due to the fact that the proposed copilot involves various tasks such as emotion recognition, diagnosis, and treatment plan design, it is sufficient to embed demographic data from different cases under various emotional states, along with the corresponding compressed EEG signals, into the training data. In downstream applications, participant data can be input into the fine-tuned model in a specific format to obtain diagnostic information and treatment plans, as shown in Fig. 7.
Figure 7: Flowchart of assisted electronic medical record generation implemented in EEG Emotion Copilot.
[…]
IV EXPERIMENTS
IV-A Implementation Details
Dataset: We utilize the FACED dataset [29] to validate our copilot. This dataset includes recordings from 123 subjects, each with 32 EEG channels, as they watched 28 video clips, each lasting 30 seconds. These clips encompass nine emotions: anger, disgust, fear, sadness, neutral, amusement, Inspiration, Joy, and Tenderness. Each positive and negative emotion is represented by three clips, while the neutral emotion is depicted in four clips.
Training Details: In the model pruning phase, we set the pruning ratio to 0.5 and then trained on the public datasets with a learning rate of 1e-5, incorporating L1 and L2 regularization terms. The training was conducted for 40 Epoches. During the LoRA fine-tuning phase, we maintained a learning rate of 1e-5, applied weight decay and gradient decay, and set LoRAr to 8, LoRAα to 32, and LoRAdropout to 0.5. The fine-tuning process spanned 15 epochs, dividing into 5 steps. Model pruning was executed on a NVIDIA A800 80G, while the fine-tuning were completed on NVIDIA GeForce RTX 4090 GPUs.
IV-B Performance Validation of Varied EEG Signal Length
Regarding the effect of the compressed signal length, we first extracted the EEG signal in the FACED data and the corresponding demography according to the Data Structure in Fig. 1. We implemented two compression methods and an original control group. The question-answering process and the corresponding results are shown in Fig. 8. Although the euclidean distance and cosine similarity of Ans1 and Ans2 are similar to the baseline for the same question, Ans2 is more consistent with the baseline than Ans1, as Ans1 differs from the baseline in terms of emotion recognition. Therefore, with limited resources, it is often more effective to compress the signal to a fixed length using wavelet transforms rather than a segmented approach.
TABLE I: Performance Comparison Of Competing Light-Weight Models
Compression Model Parameters Storage_size Precision Epoch Nine Emotion Recognition Three Emotion Recognition
Method F1 Avg. RT (s) F1 Avg. RT (s)
[…]
checkpoint1-1 0.15B 0.294G bf16 3 0.080 0.430 0.216 0.420
checkpoint1-2 0.15B 0.294G bf16 3 0.096 0.510 0.324 0.510
checkpoint1-3 0.15B 0.294G bf16 3 0.100 0.540 0.376 0.540
checkpoint1-4 0.15B 0.294G bf16 3 0.093 0.600 0.359 0.610
checkpoint1-5 0.15B 0.294G bf16 3 0.103 0.610 0.340 0.580
base_model-1 [22] 0.50B 0.920G bf16 6 0.143 1.120 0.445 1.220
opt-350m-1 [58] 0.35B 0.647G bf16 15 0.036 3.240 0.037 3.180
LiteLlama 0.46B 0.901G bf16 15 0.041 1.910 0.057 1.900
gpt-sw3-356m 0.47B 1.580G bf16 15 0.035 1.160 0.035 1.180
W Our Model 2
checkpoint2-1 0.15B 0.294G bf16 3 0.118 0.580 0.432 0.590
checkpoint2-2 0.15B 0.294G bf16 3 0.346 0.590 0.980 0.590
checkpoint2-3 0.15B 0.294G bf16 3 0.339 0.550 0.960 0.540
base_model-2 [22] 0.50B 0.920G bf16 6 0.153 1.230 0.450 1.290
opt-350m [58] 0.35B 0.647G bf16 15 0.037 3.960 0.038 4.090
LiteLlama 0.46B 0.901G bf16 15 0.036 2.120 0.036 2.120
gpt-sw3-356m 0.47B 1.580G bf16 15 0.036 1.070 0.036 1.260
[…]
IV-C Performance Comparison with Lightweight Models
To better validate our experimental hypothesis, we combine the strategies outlined in Fig. 6 and Fig. 8. First, we divide the training process based on the data derived from compression method 1, into five steps, with each step trained for 3 epochs. For each step’s checkpoint (i.e., checkpoint1-n), we perform classification of content generation for the nine emotions described in III-A, followed by classification for three emotions (positive, negative, neutral). For comparison, we use Qwen2-0.5B as the base model, training it for 6 epochs under identical experimental conditions to obtain base_model-1, and perform classification, as shown in Table I.
Next, we fine-tuned the models based on the data derived from compression method 2. Specifically, we fine-tuned the pruned model in three steps, with 3, 3, and 3 epochs respectively, and perform classification for both the nine emotions and the three emotions. For consistency, Qwen2-0.5B is also trained for 6 epochs to obtain base_model-2, after which the corresponding classification experiment is conducted, as shown in Table I.
Additionally, we include opt-350m [58], LiteLlama, and gpt-sw3-356m as baseline references. Indeed, the checkpoint1-n model of Model 1 did not achieve ideal results on both tasks compared to the base_model-1. However, considering the F1 scores across multiple epochs of training, the model’s performance on this task gradually improved. In Model 2, the performance of checkpoint2-2 and checkpoint2-3 showed significant improvement compared to models with the similar parameter size, and they also consumed less average response time.
IV-D Performance Comparison with Heavyweight Models
In Table II, we present the results of emotion recognition for Qwen2-1.5B [22], InternLM2.5-1.8B [59], Qwen2.5-3B [60], and Qwen2.5-7B [60], and InternLM2.5-7B [59]. Surprisingly, the results were not as impressive as expected. The reason for this is that the default setting for top-k is 50. But even if top-k is set to 1, the results are still not ideal. For the heavyweight models, their responses are more flexible due to their larger knowledge storage, allowing them to generate a wider range of answers. This also indicates that, despite the very low loss during training, which might suggest model fitting, the actual responses in specific scenarios are dissatisfactory.
Figure 9: t-SNE [61] visualization of K-Means (A-B) and spectral clustering (C-D) results of embeddings based on the EEG signal length=500 (compression method 1) and EEG signal length=50 (compression method 2) in Fig. 8, respectively.
Figure 10: Clustering results of data for compression method 1 and 2 described in Fig. 8 using various models mentioned in Table I and II with K-Means clustering and t-SNE [61] visualization of the text embeddings.
Furthermore, we first use the pre-trained all-MiniLM-L6-v2 model for text representation. Specifically, we combine the “prompt” and “treatment” from each record into a single string and generate the corresponding embedding vectors. These embeddings are then subjected to K-Means clustering and spectral clustering [62] for unsupervised learning, with the number of clusters set to 9. The t-SNE [61] is applied to reduce the dimensionality of the embedding and visualize the results, as shown in Fig. 9.
In addition, we investigate the clustering performance for data of compression method 1 and 2 respectively in Fig. 8 by applying the same procedure to a range of models included in Table I and II. We convert the text into the appropriate input format using the tokenizer, extract the hidden states from the final layer, and apply average pooling to generate fixed-length text embedding vectors. These embeddings are clustered using K-Means, and dimensionality reduction is performed using the t-SNE for visualization, as shown in Fig. 10. Indeed, as illustrated in Fig. 9 and 10, the visualization suggests that corpora of two different lengths are not easily clustered. Specifically, compared to the data of Model 1, the data of Model 2 exhibits a higher degree of aggregation, which directly corresponds to a reduction in computational load.
Figure 11: Performance comparison as to F1 and Avg. RT of nine and three emotion classification across steps of two fine-tuning strategies.
IV-E Model Fine-tuning Strategy Comparison After Pruning
To evaluate the two fine-tuning strategies outlined in Fig. 5, we adopted compression method 2 show in Fig. 8 for comparison. According to the results from Table I and Fig. 11, we found that Strategy 1 demonstrated better performance in Table I than Strategy 2 did in the lightweight models fine-tuning process.
TABLE II: Performance comparison Of Heavy-Weight Models
Compression Model Parameters Storage Size Precision Epoch Nine Emotion Recognition Three Emotion Recognition
Method F1 (Top-K=50) Avg. RT (s) F1 (Top-K=50) Avg. RT (s)
[…]
Internlm2.5-1.8b-1 [59] 1.8B 3.780G bf16 3 0.036 5.870 0.035 5.910
qwen2.5-3b-1 [60] 3B 6.170G bf16 3 0.111 4.260 0.041 10.960
qwen2.5-7b-1 [60] 7B 15.230G bf16 2 0.051 1.410 0.152 2.100
Internlm2.5-7b-1 [59] 7B 14.200G bf16 2 0.082 10.090 0.039 10.090
W qwen2-1.5b-2 [22] 1.5B 3.091G bf16 3 0.074 3.730 0.328 3.510
Internlm2.5-1.8b-2 [59] 1.8B 3.780G bf16 3 0.036 6.320 0.041 4.470
qwen2.5-3b-2 [60] 3B 6.170G bf16 3 0.133 3.680 0.114 3.760
qwen2.5-7b-2 [60] 7B 15.230G bf16 2 0.036 0.020 0.038 0.240
Internlm2.5-7b-2 [59] 7B 14.200G bf16 2 0.082 10.260 0.042 7.300
Compression Model Parameters Storage Size Precision Epoch Nine Emotion Recognition Three Emotion Recognition
Method F1 (Top-K=1) Avg. RT (s) F1 (Top-K=1) Avg. RT (s)
[…]
Internlm2.5-1.8b-1 [59] 1.8B 3.780G bf16 3 0.036 7.100 0.036 5.830
qwen2.5-3b-1 [60] 3B 6.170G bf16 3 0.080 3.810 0.342 4.500
qwen2.5-7b-1 [60] 7B 15.230G bf16 2 0.040 0.690 0.036 0.870
Internlm2.5-7b-1 [59] 7B 14.200G bf16 2 0.090 8.950 0.040 9.390
W qwen2-1.5b-2 [22] 1.5B 3.091G bf16 3 0.081 3.780 0.292 4.170
Internlm2.5-1.8b-2 [59] 1.8B 3.780G bf16 3 0.036 7.210 0.036 4.470
qwen2.5-3b-2 [60] 3B 6.170G bf16 3 0.036 6.320 0.038 4.470
qwen2.5-7b-2 [60] 7B 15.230G bf16 2 0.051 4.870 0.110 4.920
Internlm2.5-7b-2 [59] 7B 14.200G bf16 2 0.081 11.640 0.041 9.790
IV-F Assisted Electronic Medical Record Generation with Our Copilot
Fig. 12 illustrates an example of using our EEG Emotion Copilot to create an electronic medical record. Initially, the starting sentence of the prompt is provided, followed by the addition of demographic information of the subjects, including age, gender, and certain facial features. Finally, the preprocessed and compressed signal is incorporated to complete the prompt, which is then input into EEG Emotion Copilot. The EEG Emotion Copilot first analyzes the data to generate an answer, which is subsequently transformed into the record through format conversion.
V Discussion
V-A Data Structure of Prompt
We developed a comprehensive human-based dataset that integrates both public knowledge datasets and specialized emotion datasets, facilitating the construction of a robust data structure to drive the development of LLMs. By employing effective prompting strategies, we enhanced the model’s ability to generate contextually relevant responses. Furthermore, while the performance of the models may not be exceptionally high, as illustrated in Table I and II, language remains more challenging to cluster unsupervised compared to other data types [63]. Nonetheless, by utilizing a pretrained sentence transformer model of an appropriate size, the model is still able to capture the underlying emotional features of the language to some extent as demonstrated in the spectral clustering shown in Fig. 9, despite potential issues with feature imbalance.
V-B Data Redundancy in EEG Signal Processing
We addressed the issue of data redundancy inherent in EEG signal processing. Through efficient data compression techniques, we optimized the handling of long sequences of EEG data, significantly improving computational efficiency and facilitating real-time emotion computation. However, simply compressing signal data is computationally intensive in many cases, and in complex scenarios, more channel signals are needed. Therefore, if LLMs can generate dense channel signals using only a few channel data and then analyze the emotions of the subjects [64], [65], computational efficiency will be greatly improved.
V-C Patient Privacy and Model Efficiency
We emphasized the importance of patient privacy by ensuring that our model can run locally. We explored model pruning strategies to create a lightweight version of the language model, making it feasible to deploy in environments with limited computational resources while maintaining performance.
V-D Performance of Lightweight Models
We performed model pruning and compared the performance of the models obtained through the two fine-tuning methods illustrated in Fig. 5. It was found that the results obtained using Strategy 1 for fine-tuning were superior. Furthermore, as shown in Table I and II, the lightweight pruned model achieved through multi-step fine-tuning demonstrated better performance compared to other models with larger parameter sizes.
V-E Hallucination of Heavyweight Models
While a low training loss suggests a good fit to the training data, it does not necessarily reflect the model’s true performance, as the generated responses may still fail to meet expectations or even diverge significantly [66]. This issue is particularly evident in 7B and larger models, where the hallucination problem becomes more pronounced in specialized domains. These findings also suggest that lightweight models may be better suited for such domain-specific generation tasks.
Figure 12: An example using our EEG Emotion Copilot to create the assisted electronic medical record. Initially, the prompt is provided as the starting sentence, followed by the addition of the subject’s demographic information, including gender, age, and certain facial features (female, 23 years old, negative). Finally, the preprocessed and compressed EEG signal is incorporated to complete the prompt, which is then input into EEG Emotion Copilot.
VI Conclusion
In this paper, we present EEG Emotion Copilot, a lightweight and locally-running LLM designed for emotional recognition using EEG signals and the generation of corresponding treatment plans for assisted electronic medical records. Our findings indicate that EEG Emotion Copilot has the potential to revolutionize the recognition and treatment of emotions in clinical settings, ultimately enhancing human-computer interaction and improving patient outcomes. Future work will combine lightweight LLMs with multimodal data such as human facial expressions, speech, gestures to achieve greater accuracy in affective computing.